BROADENING AND SHIFT COEFFICIENTS IN THE 3-μm BANDS OF HCN

\underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; and M. DULICK, National Solar Observatory, P. O. Box 26732, Tucson, AZ 85726-6732.

While there have been a number of recent studies of room-temperature N_{2}-broadening of spectral lines in various molecular bands of HCN, the corresponding pressure-induced line shifts have not been measured. Only three limited studies of the temperature-dependence of N_{2}-broadening in HCN have been reported, and air- and O_{2}-broadening have been examined only at room temperature for single HC^{15}N rotational line.^{a} We have recorded over 80 new infrared spectra of HCN broadened by N_{2}, O_{2}, or air at room temperature and low temperatures using a 50-cm coolable cell with the McMath-Pierce Fourier transform spectrometer (FTS) of the National Solar Observatory on Kitt Peak, Arizona. Most of these spectra cover bandpasses of either 2750-3950 cm^{−1} at 0.008 cm^{−1} resolution or 600-3000 cm^{−1} at 0.005 cm^{−1} resolution. Volume mixing ratios of HCN in the broadening gases were 0.1% or less, and total sample pressures ranged from 12 Torr to over 500 Torr. Most spectra were recorded at temperatures from room temperature down to near the freezing point of HCN (−13^{∘}C), and some spectra were successfully recorded at temperatures down to about −60^{∘}C. Several spectra of low pressures (about 0.1 Torr) of purified HCN were also recorded to provide line positions unaffected by pressure-induced shifts. An initial group of 19 of these absorption spectra have been analyzed simultaneously using our multispectrum nonlinear least-squares technique^{b} to determine N_{2}-broadening and shift coefficients and their temperature dependences for transitions up to J′′ = 30 in the ν_{1} fundamental band of H^{12}C^{14}N. The results for N_{2}-broadening compare well with available values reported in the literature. In particular, the N_{2}-broadening temperature-dependence exponent n shows a quantum-number dependence similar to that observed in the ν_{2} band by Schmidt et al.^{c} Analysis of air-broadening and shifts in the ν_{1} band is in progress, as well as work on the ν_{2} and 2ν_{2} band systems.

aJ.-M. Colmont, J. Mol. Spectrosc. \underline{114}, 298-304 (1985).

bD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

cC. Schmidt, J.-C. Populaire, J. Walrand, G. Blanquet and J.-P. Bouanich, J. Mol. Spectrosc. \underline{158}, 423-432 (1993).